OpenStudy (anonymous):

integrate (1/2t) - (sqrt 2)(e^t) dt

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 1 }{ 2t }-\sqrt{2}e^{2}dt\]

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 1 }{ 2t } dt - \int\limits_{}^{} \sqrt{2}e^2 dt\]

OpenStudy (anonymous):

@hartnn help :)

hartnn (hartnn):

constants can be taken out of integral \(\large \int 1/t dt = \ln t +c \\ \large \int e^t dt = e^t +c \)

OpenStudy (anonymous):

@hartnn  i dont get the last one. its e^2 dt. how? sorry for the trouble. >.<